Mary's family is driving to Denver, which is 560 miles away. If they average 70 mph, and take 30 minutes breaks every 2 hours, how long will it take?
Answer by Cromlix(4381)

Hi there,
Time = Distance/Speed
Time = 560/70 = 8 hours
8 hours = 4 x 2 hours
4 x 30 minutes = 2 hours
Mary's full journey to Denver will
take 8 + 2 hours = 10 hours total.
Hope this helps :-)